# spacing=nonfrench

### 20 Mar 2026 v1

@@ Abstract
Formal verification via interactive theorem proving is increasingly used to ensure the correctness of critical systems, yet constructing large proof scripts remains highly manual and limits scalability. Advances in large language models (LLMs), especially in mathematical reasoning, make their integration into software verification increasingly promising. This paper introduces a neuro-symbolic proof generation framework designed to automate proof search for systems-level verification projects. The framework performs a best-first tree search over proof states, repeatedly querying an LLM for the next candidate proof step. On the neural side, we fine-tune LLMs using datasets of proof state–step pairs; on the symbolic side, we incorporate a range of ITP tools to repair rejected steps, filter and rank proof states, and automatically discharge subgoals when search progress stalls. This synergy enables data-efficient LLM adaptation and semantics-informed pruning of the search space. We implement the framework on a new Isabelle REPL that exposes fine-grained proof states and automation tools, and evaluate it on the FVEL seL4 benchmark and additional Isabelle developments. On seL4, the system proves up to 77.6% of the theorems, substantially surpassing previous LLM-based approaches and standalone Sledgehammer, while solving significantly more multi-step proofs. Results across further benchmarks demonstrate strong generalization, indicating a viable path toward scalable automated software verification.
@@ 1 Introduction
Formal verification plays a vital role in ensuring the correctness and reliability of software systems, especially in safety- and security-critical domains where failures can be catastrophic (Clarke1996, ; Blanchet2003, ; Woodcock2009, ). Interactive theorem proving (ITP) stands out for its expressiveness and ability to provide the strongest assurance. With ITPs, landmark projects such as CompCert (a formally verified optimizing C compiler whose correctness is mechanically proven in Rocq leroy2016compcert ) and seL4 (formally verified OS microkernel whose functional correctness, security properties and reliability guarantees are proven in Isabelle/HOL klein2014comprehensive ) have been successfully delivered, which demonstrate that formally verified systems can achieve both practical usability and rigorous correctness.
While ITP offers unparalleled assurance, it also incurs substantial costs, preventing its wide adoption, especially in large-scale, systems-level software (ringer2019qed, ). Specifically, a typical software verification workflow involves 1) writing precise formal specifications to be proved and 2) constructing rigorous proofs for these specifications, both of which demand immense human effort. For example, the seL4 verification project required approximately 20 person-years of work, and produced over 100K lines of proof scripts, compared to only about 10K lines of C implementation and merely 3K lines in its simplest abstract specification in Isabelle klein2014comprehensive . Moreover, developing such specifications and proofs necessitates specialized expertise in both theorem proving and the target domain, a combination that is challenging to acquire.
In this work, we focus on the proof generation step in systems-level software verification projects. Recently, large language models (LLMs) have demonstrated encouraging mathematical reasoning capabilities, which inspire research into their application for software verification, particularly for proof generation  qin2025can ; thompson2024rango ; zhang2024selene ; lin2024fvel . Unfortunately, both prior studies and our initial experiments show that, even when theorem statements are well formalized and relevant libraries are fully provided in advance, current LLMs still struggle to generate complete, correct proofs. We summarize two challenges as follows.
@@
❷ Pausity of usable data. The second challenge, which is closely related to the first, is the shortage of high-quality (training) data for software verification. For instance, the seL4 project contains only around 20K theorems and lemmas, and most proofs are written in a procedural style, where much of the reasoning is implicit and encoded in the interaction with the theorem prover. This substantially limits their usefulness in effective LLM training/fine-tuning. Indeed, FVEL lin2024fvel , which systematically extracted theorems from seL4 to fine-tune Mistral-7B-Instruct jiang2023mistral and Llama-3-8B-Instruct dubey2024llama models, reported a success rate below 10% (88 out of 1,077 theorems in the test set). [^1] Selene only extracts a small subset of the seL4 project, and its collected theorems are relatively easy, with proof lengths ranging from one to five. In contrast, the FVEL benchmark is more comprehensive and challenging, as it extracts all the theorems to be proved from seL4.
@@
Approach. In this paper, we propose a proof generation framework that synergistically combines LLMs and ITPs, aiming for automated systems-level software verification. Our approach adopts a proof step-based tree-search pipeline that iteratively queries a fine-tuned LLM to predict the next proof step until the target theorem is successfully proved.
@@
To address the first challenge, our framework adopts a neuro-symbolic approach tailored to the nature of software verification. First, to address challenges posed by domain-specific lemmas and tactics, we incorporate a proof-step revision module that repairs or refines LLM-generated steps whenever they fail to compile or make no progress. Second, to efficiently prune unpromising branches, we utilize Nitpick (blanchette2010nitpick, ) and QuickCheck (claessen2000quickcheck, ) to “test” each candidate proof state and discard those that admit counterexamples; the remaining candidates are then ranked by the LLM according to their plausibility. Finally, if the tree search still fails to resolve the goal, Sledgehammer (bohme2010sledgehammer, ) acts as a backstop that scans the local library for potentially relevant lemmas and attempts to complete the current proof.
@@
To address the second challenge, our tree-search pipeline facilitates direct interaction between the LLM and the theorem prover, by extracting the prover’s internal proof state to LLM, which can inform and guide subsequent step predictions. This pipeline also enables “bootstrapping” training data generation. Namely, during the tree-search process, training data for proof-step prediction can be extracted on-the-fly. For example, if a theorem’s proof consists of five steps, a successful tree search could yield five training instances, which can later be used for, e.g., fine-tuning (wu2024internlm2, ; xin2025bfs, ).
@@
Experiments. We evaluate our framework on seL4 klein2014comprehensive proof construction. The results show that, with a fine-tuned language model, our approach achieves a total proof success rate of up to 77.6%, significantly surpassing prior automated proof-generation approaches for seL4. Noteably, it excels, compared to the two preceding systems, at successfully handling a larger number of multi-step proofs, a challenge that has not been addressed by prior work. Moreover, our framework substantially alleviates manual proof-engineering effort, reducing expert proof-writing effort by 71.1% on average in an AI–human collaboration setting. Finally, experimental results on benchmarks from Archive of Formal Proofs projects (e.g., X86 Semantics , IEEE Floating Point, and SATSolverVerification), as well as the code verification benchmark Code2Inv (by translating its loop invariants into Isabelle theorems) confirm the strong generalizability of our approach.
@@
Summary of main contributions. (1) We provide a neuro-symbolic framework based on LLM-enabled tree search, tailored for real-world, systems-level software verification projects. (2) We implement concrete integration of LLMs and a broad range of symbolic, logic-based tools, substantially enhancing the efficacy of proof-tree search; (3) We curate training datasets and benchmarks, and carry out thorough empirical evaluations. Our work advances a new frontier of AI4Verification, an emerging area focused on applying cutting-edge AI techniques to hardware/software verification, especially to support large-scale proof engineering for safety-critical systems and architectures.
@@
Figure 1: An example theorem and its human-written proof from the seL4 project, shown alongside LLM-generated attempts. The human-written Isabelle proof illustrates the procedural, tactic-driven reasoning style that is common in software verification. We prompted two state-of-the-art LLMs, Gemini 3 and GPT-5.1, to synthesize the same proof under controlled conditions, with web search disabled to avoid any leakage of the original script. Although both models correctly recognized the proof context and invoked some relevant tactics, neither succeeded in producing a valid proof even after roughly one minute of internal reasoning. Their attempts frequently misuse the domain-specific wp tactic, or hallucinate nonexistent lemmas.
@@
Structure. Section 2 illustrates the challenges of the problem. Section 3 presents our neuro-symbolic approach. Section 4 reports the evaluation of our approach on seL4 and four additional repository-level benchmarks. Section 5 reviews the related work. Section 6 discusses the limitations of the current work and outlines directions for future research. Section 7 concludes the paper.
@@ 2 Problem
To show why existing LLMs struggle to synthesize proofs in real-world software verification tasks, we present a representative theorem from seL4. We choose seL4 as it remains the most comprehensive and influential milestone for system software verification. Its entire C implementation is proven functionally correct against a formal specification in Isabelle/HOL, ensuring the absence of crashes and undefined behavior. Building on this foundation, seL4’s verification was extended to enforce noninterference of information flows and access control, establishing formal guarantees of confidentiality and integrity. To date, seL4 remains one of the safest OS kernels with machine-checked proofs of both implementation-level correctness and security properties, setting the benchmark for formal system verification.
@@
To assess LLMs in generating this proof automatically, we experiment with two state-of-the-art proprietary models, i.e., Gemini 3 Pro (google2024gemini, ) and GPT-5.1 (openai2024gpt4techreport, ); their results are shown on the right panel of the figure. While both deep-thinking models successfully recognize the seL4 proof context, they are unable to apply the domain-specific tactic wp correctly. They also hallucinate lemmas that do not exist in the library (e.g., hoare_seq_ext). We further attempted several rounds of iterative repair by feeding back the ITP’s error messages, but the regenerated proofs continued to fail verification. An analysis of the models’ reasoning traces shows that they frequently misinterpret the functionality and performance characteristics of the wp tactic and related lemmas (such as set_cap_valid_objs).
There are two potential approaches to addressing this issue. The first is to use retrieval-augmented generation (yang2023leandojo, ; thompson2024rango, ). In this paradigm, given a target theorem, a retrieval module searches for relevant or structurally similar theorems and proofs, which are then supplied to the LLM as contextual guidance. However, this strategy depends on both a high-quality proof corpus and an effective retrieval model: requirements that are difficult to satisfy in our setting. For example, in seL4, when developers introduce a new module (such as SysInit, as shown in our evaluation) together with its accompanying proofs, comparable proofs are often absent from existing libraries. Moreover, even when analogous theorems do exist, neither model-agnostic nor model-dependent retrieval techniques reliably identify them within a large verification codebase (berlot2025llm, ; sesterhenn2025compute, ). In such cases, weak retrieval can degrade, rather than improve, the quality of generated proofs.
@@ 3 Methodology
We propose a neuro-symbolic approach for proof step search on systems-level theorems. The overall procedure is illustrated in Figure 2. Starting from the root node, which represents the initial proof state, the search alternates between proposing new steps and deriving new proof states by executing those steps within the ITP.
To construct the next set of candidate proof steps for the current proof state, we first prompt a fine-tuned LLM to generate potential steps, and then apply a rule-based revision to refine or fix any imprecise suggestions. After executing these steps and obtaining a collection of new proof states, we introduce a hybrid mechanism to identify the most promising ones for further exploration. Specifically, we apply QuickCheck (claessen2000quickcheck, ) and Nitpick (blanchette2010nitpick, ), which serve as counterexample generators that help detect false conjectures or invalid intermediate proof states in Isabelle/HOL, to discard any proof state that contains a counterexample. We then use the LLM’s prediction score (cumulative perplexity) to rank the remaining states and select the most promising one. Finally, if the search for the proof fails, we invoke an automated tool, i.e., Sledgehammer, to attempt to complete the proof. Putting all the above together, we establish the framework in Algorithm 1.
@@ 3.1.1 Proof Step Generation
The rationale for focusing on the current proof state is twofold. First, automated theorem proving can be cast as a sequential decision-making problem, and thus can be modeled as a Markov decision process, in which a standard result stipulates that pure, history-independent polices (i.e., the next step depends solely on the present state) suffice for the reachability objective, i.e., to reach the “proof done” state. This is also a common practice in the literature of proof search, which turns out to be critical for efficiency (wu2021tacticzero, ; huang2025leanprogress, ; wu2020neural, ). Second, the structure and vocabulary of the proof state provide strong cues regarding which tactics are applicable and which premises are likely to be relevant (gauthier2021tactictoe, ; piotrowski2023machine, ). Additionally, we focus on predicting a single step at a time. This is because, although multi-step prediction is possible in principle, the complexity of real proofs and the limited amount of available data make such predictions considerably more challenging; empirically, one-step prediction yields more reliable results.
@@ 3.1.2 Proof Step Revision
5:for each $(s,\delta,\log p_{\delta})\in$ failStats do
@@ 3.2.1 Proof State Filtering
To alleviate state explosion in tree-search approaches, we reorganize the accumulated proof states before further exploration. Particularly, after generating new proof states by executing the newly produced proof steps, we first invoke two complementary symbolic tools, QuickCheck (claessen2000quickcheck, ) and Nitpick (blanchette2010nitpick, ), to detect potential counterexamples in the candidate states. QuickCheck performs random property-based testing on executable parts of the proof state, whereas Nitpick translates higher-order formulas into finite relational models and searches for counterexamples using Kodkod, which is a SAT-based first-order relational model finder.
Additionally, we adapt the tool SolveDirect (SolveDirect2025, ) to detect potentially duplicate proof states. Initially, SolveDirect determines whether a newly stated theorem can be solved directly using an existing one. In our setting, we formalize two proof states as theorems and apply SolveDirect in both directions to determine whether they are semantically equivalent.
@@
theorem sofl_test:
sint x + sint y = sint (x + y) $\Longleftrightarrow$
drop_bit (size x - 1)
((x + y XOR x) AND (x + y XOR y)) = 0
for x y :: a::len word

### 07 May 2026 v2

@@ Abstract
Formal verification via interactive theorem proving is increasingly used to ensure the correctness of critical systems, yet constructing large proof scripts remains highly manual and limits scalability. Advances in large language models (LLMs), especially in mathematical reasoning, make their integration into software verification increasingly promising. This paper introduces a neuro-symbolic proof generation framework designed to automate proof search for system-level verification projects. The framework performs a best-first tree search over proof states, repeatedly querying an LLM for the next candidate proof step. On the neural side, we fine-tune LLMs using datasets of proof state–step pairs; on the symbolic side, we incorporate a range of ITP tools to repair rejected steps, filter and rank proof states, and automatically discharge subgoals when search progress stalls. This synergy enables data-efficient LLM adaptation and semantics-informed pruning of the search space. We implement the framework on a new Isabelle REPL that exposes fine-grained proof states and automation tools, and evaluate it on the FVEL seL4 benchmark and additional Isabelle developments. On seL4, the system proves up to 77.6% of the theorems, substantially surpassing previous LLM-based approaches and standalone Sledgehammer, while solving significantly more multi-step proofs. Results across further Isabelle benchmarks demonstrate strong generalization, indicating a viable path toward scalable automated software verification.
@@ 1 Introduction
Formal verification plays a vital role in ensuring the correctness and reliability of software systems, especially in safety- and security-critical domains where failures can be catastrophic (Clarke1996; Blanchet2003; Woodcock2009). Interactive theorem proving (ITP) stands out for its expressiveness and ability to provide the strongest assurance. With ITPs, landmark projects such as CompCert (a formally verified optimizing C compiler whose correctness is mechanically proven in Rocq leroy2016compcert) and seL4 (a formally verified OS microkernel whose functional correctness, security properties and reliability guarantees are proven in Isabelle/HOL klein2014comprehensive) have been successfully delivered, which demonstrate that formally verified systems can achieve both practical usability and rigorous correctness.
While ITP offers unparalleled assurance, it also incurs substantial costs, preventing its wide adoption, especially in large-scale, system-level software (ringer2019qed). Specifically, a typical software verification workflow involves 1) writing precise formal specifications to be proved and 2) constructing rigorous proofs for these specifications, both of which demand immense human effort. For example, the seL4 verification project required approximately 20 person-years of work, and produced over 100K lines of proof scripts, compared to only about 10K lines of C implementation and merely 3K lines in its simplest abstract specification in Isabelle klein2014comprehensive. Moreover, developing such specifications and proofs necessitates specialized expertise in both theorem proving and the target domain, a combination that is challenging to acquire.
In this work, we focus on the proof generation step in system-level software verification projects. Recently, large language models (LLMs) have demonstrated encouraging mathematical reasoning capabilities, which inspire research into their application for software verification, particularly for proof generation  qin2025can; thompson2024rango; zhang2024selene; lin2024fvel. Unfortunately, both prior studies and our initial experiments show that, even when theorem statements are well formalized and relevant libraries are fully provided in advance, current LLMs still struggle to generate complete, correct proofs. We summarize two challenges as follows.
@@
❷ Paucity of usable data. The second challenge, which is closely related to the first, is the shortage of high-quality (training) data for software verification. For instance, the seL4 project contains only around 20K theorems and lemmas, and most proofs are written in a procedural style, where much of the reasoning is implicit and encoded in the interaction with the theorem prover. This substantially limits their usefulness in effective LLM training/fine-tuning. Among existing seL4 benchmarks, Selene zhang2024selene extracts a small subset of the project, targeting relatively easy theorems with proof lengths of one to five. FVEL lin2024fvel is more comprehensive and challenging, as it extracts all theorems to be proved from seL4. Yet even with systematic fine-tuning of Mistral-7B-Instruct jiang2023mistral and Llama-3-8B-Instruct dubey2024llama, FVEL reported a success rate below 10% (88 out of 1,077 theorems in the test set).
@@
Approach. In this paper, we propose a proof generation framework that synergistically combines LLMs and ITPs, aiming for automated system-level software verification. Our approach adopts a proof step-based tree-search pipeline that iteratively queries a fine-tuned LLM to predict the next proof step until the target theorem is successfully proved.
@@
To address the first challenge, our framework adopts a neuro-symbolic approach tailored to the nature of software verification. First, to address challenges posed by domain-specific lemmas and tactics, we incorporate a proof-step revision module that repairs or refines LLM-generated steps whenever they fail to compile or make no progress. Second, to efficiently prune unpromising branches, we utilize Nitpick (blanchette2010nitpick) and QuickCheck (claessen2000quickcheck) to “test” each candidate proof state and discard those that admit counterexamples; the remaining candidates are then ranked by the LLM according to their plausibility. Finally, if the tree search still fails to resolve the goal, Sledgehammer (bohme2010sledgehammer) acts as a backstop that scans the local library for potentially relevant lemmas and attempts to complete the current proof.
@@
To address the second challenge, we construct our training data by extracting internal proof states from existing theorems. For example, a five-step proof yields five distinct training instances, which are subsequently used for fine-tuning. During inference, our tree-search pipeline leverages this state extraction to facilitate direct interaction between the LLM and the prover, informing step predictions. While we currently train on existing human-authored proofs, this pipeline also enables future “bootstrapping”, where new training data could be extracted on-the-fly from successful search paths (wu2024internlm2; xin2025bfs).
@@
Experiments. We evaluate our framework on seL4 klein2014comprehensive proof construction. The results show that, with a fine-tuned language model, our approach achieves a total proof success rate of up to 77.6%, significantly surpassing prior automated proof-generation approaches for seL4. Notably, it excels, compared to the two preceding systems, at successfully handling a larger number of multi-step proofs, a challenge that has not been addressed by prior work. Moreover, our framework substantially alleviates manual proof-engineering effort, reducing expert proof-writing effort by 71.1% on average in an AI–human collaboration setting. Finally, experimental results on benchmarks from Archive of Formal Proofs projects (e.g., X86 Semantics , IEEE Floating Point, and SATSolverVerification), as well as the code verification benchmark Code2Inv (by translating its loop invariants into Isabelle theorems) confirm the strong generalizability of our approach.
@@
Summary of main contributions. (1) We provide a neuro-symbolic framework based on LLM-enabled tree search, tailored for real-world, system-level software verification projects. (2) We implement concrete integration of LLMs and a broad range of symbolic, logic-based tools, substantially enhancing the efficacy of proof-tree search; (3) We curate training datasets and benchmarks, and carry out thorough empirical evaluations. Our work advances a new frontier of AI4Verification, an emerging area focused on applying cutting-edge AI techniques to hardware/software verification, especially to support large-scale proof engineering for safety-critical systems and architectures.
@@
Figure 1: An example theorem and its human-written proof from the seL4 project, shown alongside LLM-generated attempts. The human-written Isabelle proof illustrates the procedural, tactic-driven reasoning style that is common in software verification. We prompted two state-of-the-art LLMs, Gemini 3 Pro and GPT-5.1, to synthesize the same proof under controlled conditions, with web search disabled to avoid any leakage of the original script. Although both models correctly recognized the proof context and invoked some relevant tactics, neither succeeded in producing a valid proof even after roughly one minute of internal reasoning. Their attempts frequently misuse the domain-specific wp tactic, or hallucinate nonexistent lemmas.
@@
Structure. Section 2 illustrates the challenges of the problem. Section 3 presents our neuro-symbolic approach. Section 4 reports the evaluation of our approach on seL4 and four additional repository-level benchmarks. Section 5 provides further qualitative analysis through a case study and an examination of failure modes. Section 6 reviews the related work. Section 7 discusses the limitations of the current work and outlines directions for future research. Section 8 concludes the paper.
@@ 2 Problem
To demonstrate why existing LLMs struggle to synthesize proofs in real-world software verification tasks, we present a representative theorem from seL4. We choose seL4 as it is one of the most widely recognized formally verified systems, with its entire C implementation proven functionally correct against a formal specification in Isabelle/HOL, ensuring the absence of crashes and undefined behavior. Building on this foundation, seL4’s verification was extended to enforce noninterference of information flows and access control, establishing formal guarantees of confidentiality and integrity. To date, seL4 remains one of the safest OS kernels with machine-checked proofs of both implementation-level correctness and security properties, setting the benchmark for formal system verification.
@@
To assess LLMs in generating this proof automatically, we experiment with two state-of-the-art proprietary models, i.e., Gemini 3 Pro (google2024gemini) and GPT-5.1 (openai2025gpt51); their results are shown on the right panel of the figure. While both deep-thinking models successfully recognize the seL4 proof context, they are unable to apply the domain-specific tactic wp correctly. They also hallucinate lemmas that do not exist in the library (e.g., hoare_seq_ext). We further attempted several rounds of iterative repair by feeding back the ITP’s error messages, but the regenerated proofs continued to fail verification. An analysis of the models’ reasoning traces shows that they frequently misinterpret the functionality and performance characteristics of the wp tactic and related lemmas (such as set_cap_valid_objs).
There are two potential approaches to addressing this issue. The first is to use retrieval-augmented generation (yang2023leandojo; thompson2024rango). In this paradigm, given a target theorem, a retrieval module searches for relevant or structurally similar theorems and proofs, which are then supplied to the LLM as contextual guidance. However, this strategy depends on both a high-quality proof corpus and an effective retrieval model: requirements that are difficult to satisfy in our setting. For example, in seL4, when developers introduce a new module (such as SysInit, as shown in our evaluation) together with its accompanying proofs, comparable proofs are often absent from existing libraries. Moreover, even when analogous theorems do exist, neither model-agnostic nor model-dependent retrieval techniques reliably identify them within a large verification codebase (blanchette2016hammering; yang2023leandojo). In such cases, weak retrieval can degrade, rather than improve, the quality of generated proofs.
@@ 3 Methodology
We propose a neuro-symbolic approach for proof step search on system-level theorems. The overall procedure is illustrated in Figure 2. Starting from the root node, which represents the initial proof state, the search alternates between proposing new steps and deriving new proof states by executing those steps within the ITP.
To construct the next set of candidate proof steps for the current proof state, we first prompt a fine-tuned LLM to generate potential steps, and then apply a rule-based revision to refine or fix any imprecise suggestions. After executing these steps and obtaining a collection of new proof states, we introduce a hybrid mechanism to identify the most promising ones for further exploration. Specifically, we apply QuickCheck (claessen2000quickcheck) and Nitpick (blanchette2010nitpick), which serve as counterexample generators that help detect false conjectures or invalid intermediate proof states in Isabelle/HOL, to discard any proof state that contains a counterexample. We then use the LLM’s prediction score (cumulative perplexity) to rank the remaining states and select the most promising one. Finally, if the search for the proof fails, we invoke an automated tool, i.e., Sledgehammer, to attempt to complete the proof. Putting all the above together, we establish the framework in Algorithm 1.
@@ 3.1.1 Proof Step Generation
The rationale for focusing on the current proof state is twofold. First, automated theorem proving can be cast as a sequential decision-making problem, and thus can be modeled as a Markov decision process, in which a standard result stipulates that pure, history-independent policies (i.e., the next step depends solely on the present state) suffice for the reachability objective, i.e., to reach the “proof done” state. This is also a common practice in the literature of proof search, which turns out to be critical for efficiency (wu2021tacticzero; huang2025leanprogress; wu2020neural). Second, the structure and vocabulary of the proof state provide strong cues regarding which tactics are applicable and which premises are likely to be relevant (gauthier2021tactictoe; piotrowski2023machine). Additionally, we focus on predicting a single step at a time. This is because, although multi-step prediction is possible in principle, the complexity of real proofs and the limited amount of available data make such predictions considerably more challenging; empirically, one-step prediction yields more reliable results.
@@ 3.1.2 Proof Step Revision
5:for each $(s,\delta,\log p_{\delta})\in$ failStates do
@@ 3.2.1 Proof State Filtering
To alleviate state explosion in tree-search approaches, we reorganize the accumulated proof states before further exploration. Particularly, after generating new proof states by executing the newly produced proof steps, we first invoke two complementary symbolic tools, QuickCheck (claessen2000quickcheck) and Nitpick (blanchette2010nitpick), to detect potential counterexamples in the candidate states. QuickCheck performs random property-based testing on executable parts of the proof state, whereas Nitpick translates higher-order formulas into finite relational models and searches for counterexamples using Kodkod, which is a SAT-based first-order relational model finder.
Additionally, we adapt the tool SolveDirect (SolveDirect2025) to detect potentially duplicate proof states. Initially, SolveDirect determines whether a newly stated theorem can be solved directly using an existing one. In our setting, we formalize two proof states as theorems and apply SolveDirect in both directions to determine whether they are semantically equivalent.